OpenStudy (anonymous):

Annuity and Compound Interest problem! how to solve for interest rate (NOT using tvm solver on a calculator) girl puts $300 each month into an ordinary annuity for 20 years. interest was compounded monthly. at the end of 20 years, the annuity was worth $147126. what was the annual interest rate? Im using the compound interest formula P(1+i)=A Im getting .0261496485 --> 2.6% but that is wrong, the answer is suppose to be 6.5%.. Did i make an error? Here is my work: 300(1+I)^240 = 147,126 (1+i)^240 = 147,126/300 (1+i)^240 = 490.42 240√(1+i)^240 = 240√(490.42) 1+i = 240√(490.42) i= 240√(490.42) -1 i = .0261496485 how do i solve the problem?

OpenStudy (experimentx):

do you know convolution?

OpenStudy (experimentx):

or solving first order DE?

OpenStudy (kirbykirby):

Is this not a matter of simply finding an annual effective rate?

OpenStudy (kirbykirby):

actually this is an annuity problem

OpenStudy (anonymous):

i am not familiar with convolution in mathematic terms.. and yes @kirbykirby it is an annuity problem

OpenStudy (anonymous):

help?? @Luis_Rivera @AccessDenied , @Hero .....anyone?

jimthompson5910 (jim_thompson5910):

Use the formula FV = C*[ ( (1+i)^n - 1)/i ] where in this case FV = 147126 C = 300 i = x (this is unknown and we're solving for it) n = 12*20 = 240 so this means FV = C*[ ( (1+i)^n - 1)/i ] turns into 147126 = 300*[ ( (1+x)^240 - 1)/x ] which can be rewritten as 147126x = 300( (1+x)^240 - 1) From here, I would use a graphing calculator to solve for x (since you have an exponent of 240, it's going to be tough to solve by hand)

OpenStudy (anonymous):

well i dont have to solve the entire problem b y hand.. i have to show work -- i can use a calc. lol THANK YOUU

jimthompson5910 (jim_thompson5910):

you're welcome